1st equation is 4x+3y=26
2nd equatiois y=3x
pls solve this problem. thank you.

Answer by Maths68(1474)

4x+3y=26..........(1)
y=3x..............(2)
Substitute the value of y from (2) to (1)
4x+3(3x)=26
4x+9x=26
13x=26
Divide both sides of above equation by 13
13x/13=26/13
x=2
Now substitute the value of x in (2)
y=3x
y=3(2)
y=6

(x,y)=(2,6)



Check
======
Put the value of x and y in (1)
4x+3y=26..........(1)
4(2)+3(6)=26
8+18=26
26=26